# TRACED: Transition-aware Regret Approximation with Co-learnability for Environment Design

### v4

@@ Abstract
Generalizing deep reinforcement learning agents to unseen environments remains a significant challenge. One promising solution is Unsupervised Environment Design (UED), a co‑evolutionary framework in which a teacher adaptively generates tasks with high learning potential, while a student learns a robust policy from this evolving curriculum. Existing UED methods typically measure learning potential via regret, the gap between optimal and current performance, approximated solely by value‑function loss. Building on these approaches, we introduce the transition-prediction error as an additional term in our regret approximation. To capture how training on one task affects performance on others, we further propose a lightweight metric called Co-Learnability. By combining these two measures, we present Transition‑aware Regret Approximation with Co‑learnability for Environment Design (TRACED). Empirical evaluations show that TRACED produces curricula that improve zero-shot generalization over strong baselines across multiple benchmarks. Ablation studies confirm that the transition-prediction error drives rapid complexity ramp‑up and that Co‑Learnability delivers additional gains when paired with the transition-prediction error. These results demonstrate how refined regret approximation and explicit modeling of task relationships can be leveraged for sample-efficient curriculum design in UED. https://geonwoo.me/traced/
@@ 1 Introduction
Deep reinforcement learning (RL) has achieved remarkable success in games, continuous control, and robotics (sutton1998reinforcement). Ideally, we want agents that generalize robustly to a broad range of unseen environments. However, hand‑crafting a training distribution that captures all real‑world variability is intractable, and agents often overfit even large training sets, performing poorly out‑of‑distribution (rl_generalization; rl_generalization2).
Unsupervised Environment Design (UED) tackles this by adapting the curriculum: a teacher module generates training tasks that challenge the student agent (PAIRED). A popular class of UED methods measures task difficulty by regret, the difference between the optimal return and the agent’s achieved return, and uses this metric to guide curriculum design (PAIRED; PLR; REPAIRED; ACCEL; azad2023clutr; Stabilizing; erlebach2024raccoon). Unfortunately, genuine regret requires knowing each environment’s optimal $Q^{*}$, which is infeasible in complex domains. Existing approaches, therefore, resort to coarse proxies such as Positive Value Loss (PVL) or maximum observed return (MaxMC) (rutherford2024no). In this paper, we refine regret estimation by augmenting PVL with a transition-prediction error term that captures how poorly a learned model predicts the environment’s dynamics. This combined signal provides a more faithful approximation of regret and a sharper basis for curriculum design.
We further introduce a metric called Co‑Learnability to quantify how training on one task benefits others. For instance, consider three 100‑word vocabulary tasks, Spanish, English, and Japanese, whose transfer patterns differ: because Spanish and English share many cognates, learning Spanish accelerates lexical access and boosts English accuracy (ramirez2013cross; co_ex1; co_ex2), reflecting high Co‑Learnability; in contrast, Japanese is typologically distant (Chiswick15012005), so gains from Japanese may transfer unefficiently to English, reflecting low Co‑Learnability. We present a lightweight estimator of Co‑Learnability that leverages observed changes in approximated regret, avoiding any additional modeling overhead in the UED loop.
@@
We propose TRACED (Transition-aware Regret Approximation with Co-Learnability for Environment Design), which combines refined regret and Co-Learnability to yield a task-priority landscape (Figure 1). We evaluate TRACED on procedurally generated MiniGrid (MG) (MinigridMiniworld23) and BipedalWalker (BW) (BipedalWalker). In MG, we compare against Domain Randomization (DR) (DR), PLR (PLR), ADD (Chung2024AdversarialDiffusion), and ACCEL (ACCEL) (the strongest baseline). In BW, we additionally include the state-of-the-art (SOTA) method CENIE (Jayden2024CENIE).
@@ 2.1 Underspecified Partially Observable MDPs (UPOMDPs)
We model our environments as underspecified partially observable Markov decision processes (UPOMDPs) following PAIRED. A UPOMDP is a tuple $\mathcal{M}=\langle A,O,\Theta,S,P_{0},P_{T},\mathcal{I},\mathcal{R},\gamma\rangle$, where $A$ is a finite set of actions, $S$ is a latent state space, and $O$ is an observation space. The observation function $\mathcal{I}:S\to\Delta(O)$ generates each observation $o_{t}$ given the true state $s_{t}$, the reward function $\mathcal{R}:S\times A\to\mathbb{R}$ and discount factor $\gamma\in[0,1)$ are shared across all levels. Crucially, $\Theta$ is a set of underspecified parameters that distinguish individual “levels”: for each $\theta\in\Theta$, the initial state is drawn from $P_{0}(\theta)\in\Delta(S)$ and transitions follow $P_{T}(s_{t+1}\mid s_{t},a_{t},\theta)$.
@@ 2.2 Unsupervised Environment Design (UED)
UED provides a series of levels with unknown parameters that are used to produce task environments automatically as a curriculum for the agent so as to efficiently train a single generalist policy $\pi_{\phi}$ across the entire parameter space $\Theta$ (PAIRED). Recent UED methods maximize regret of the agent to generate a distribution of environments that guide effective learning (PLR; ACCEL; Jayden2024CENIE). The regret of policy $\pi$ is defined as the difference in expected reward between optimal policy $\pi^{*}$ and policy $\pi$, $Regret_{\theta}(\pi)=U_{\theta}(\pi^{*})-U_{\theta}(\pi)$.
Because the true optimal policy $\pi^{\star}$ is unknown in complex environments, existing methods approximate the instantaneous regret using proxy metrics. For example, PLR (REPAIRED) evaluates two such proxies: Positive Value Loss (PVL) and Maximum Monte Carlo (MaxMC). PVL estimates the instantaneous regret as the average positive part of the Generalized Advantage Estimation (GAE)‐based Temporal Difference (TD) errors over an episode. PVL for an episode $\tau$ of length $T$ is defined as
@@
where $\gamma$ is the discount factor, $\lambda$ is the GAE coefficient and $\delta_{t}=r_{t}+\gamma\,V(s_{t+1})-V(s_{t})$ is the one‐step TD‐error at timestep $t$. MaxMC uses the highest undiscounted return observed on the task instead of a bootstrap target. Other criteria such as policy entropy, one‑step TD error, GAE, policy min‑margin, and policy least‑confidence have also been evaluated in curriculum learning contexts (PLR).
@@ 3.2 Task Priority Construction
A central challenge in UED is to decide which tasks to present to the agent at each step (Hughes2024OpenEndednessICML). Intuitively, we want to (1) focus on tasks that remain challenging, while (2) exploiting tasks whose training yields transfer to others. To this end, we introduce two complementary quantities, Task Difficulty and Co‑Learnability, and combine them into a single Task Priority score. In the following sections, time $t$ indexes the teacher’s curriculum-selection cycles (not environment time).
@@
Definition 4 (Task Priority). We combine difficulty and Co-Learnability into a scalar score and then apply a rank transform, which replaces raw values with their relative order (e.g., the largest value receives rank 1, the next largest rank 2, etc.), as in ACCEL (ACCEL):
@@ 3.3 Overall UED Workflow
The overall UED algorithm follows the ACCEL loop (ACCEL), with the sole change that task scoring uses our Task Priority (Eq. 7) in place of Positive Value Loss (PVL) alone (Figure 3). At each curriculum update time $t$, with probability $d$ we sample a new level uniformly at random, and with probability 1d1-d we sample a level from the replay buffer according to its Task Priority score. Early in training, when the buffer is empty, we sample levels uniformly at random for a warm-up period. The warm-up length matches ACCEL’s setting.
@@ 4 Experiments
We evaluate TRACED on two procedurally generated domains: MiniGrid (MG) and BipedalWalker (BW). In both environments, we compare against DR, PLR, ACCEL (our primary baseline, since TRACED builds directly upon it), and ADD. In BW only, we additionally include the recent sota method CENIE (Jayden2024CENIE). While SFL (rutherford2024no) is not included in the main experimental suite, we provide a comprehensive comparison against SFL in the multi-agent JaxNav environment in Appendix Q, following its original evaluation protocol. Detailed descriptions of all baselines can be found in Appendix K.
For each domain, we track the emergent complexity of the sampled curricula during training and evaluate zero‐shot test performance on held‐out levels. To summarize test results, we report the median, interquartile mean (IQM), mean, and optimality gap using the rliable library (Agarwal2021Reliable). The optimality gap measures how far an algorithm falls short of a target performance level, beyond which further gains are deemed negligible. Accordingly, higher IQM and lower optimality gap values indicate better performance. Shaded regions in all plots denote 95% confidence intervals.
All methods use Proximal Policy Optimization (PPO) (Schulman2017PPO) as the student agent. Following prior work (ACCEL; Jayden2024CENIE), we plot performance versus the number of PPO updates and report the corresponding environment interactions per update in Appendix F. Our only deviation from ACCEL is the number of PPO workers (i.e., parallel trajectory collectors), chosen for computational constraints: MiniGrid uses 16 workers instead of 32, and BipedalWalker uses 4 instead of 16. Consequently, some baseline scores may differ from those in the ACCEL paper. However, Appendix A.4 shows that TRACED has no worker-specific tuning and that reducing workers does not render the comparison unfair. We report TRACED at 10k updates (rather than 20k) because single-seed, long-horizon runs (45k PPO updates) on MiniGrid reveal post-convergence oscillations in both TRACED and ACCEL, and reporting earlier avoids this confound (Appendix B). Additional hyperparameters and architectures are provided in Appendix H.
@@ 4.1 Partially Observable Navigation
We evaluate our curriculum design on a partially observable maze navigation domain based on MiniGrid (MinigridMiniworld23), as in prior UED work. Four representatives are shown in Figure 5. The agent observes a 147‑dimensional pixel observation and is trained for up to 20k PPO updates.
@@ 4.2 Walking in Challenging Terrain
We further validate our curriculum in the continuous‐control BipedalWalkerHardcore environment from OpenAI Gym (Brockman2016Gym), as modified by POET. This domain features a procedurally generated terrain controlled by eight parameters, terrain roughness, pit gap frequency, stump height, stair spacing, etc., that jointly determine locomotion difficulty. We consider the complete set of eight parameters in our design space. Figure 5 illustrates two representative terrains with varying stair heights and surface roughness. We evaluate zero‑shot transfer on six held‑out test terrains over 100 episodes each, averaged over five random seeds.
@@ 4.3 Ablation Study
To isolate component contributions, we compare TRACED at 10k updates with two ablations (Figure 11): ATPL only (TRACED - CL) and Co-Learnability only (TRACED - ATPL), each evaluated with five seeds. Results are reported following Agarwal2021Reliable. On MiniGrid, TRACED outperforms both variants across four metrics, indicating that each component contributes meaningfully to performance. The same trend holds on BipedalWalker (Appendix A). Ablations on the scaling factors $\alpha$ and $\beta$ further show that both elements play significant roles in TRACED and that carefully balancing ATPL and CL can yield additional gains (Appendix A.1). Detailed numerical results are provided in Appendix P.
@@ 4.4 Analysis on Curriculum Progression
To analyze how the curriculum evolves, we examine the BipedalWalker level buffer over time and categorize each generated level into three difficulty bands: Easy, Moderate, and Challenging, based on environment hyperparameters. Following Jayden2024CENIE, we set thresholds for Stump Height (2.4), Pit Gap (6), Ground Roughness (4.5), and Stairs Height (5). Levels exceeding no thresholds are labeled Easy, those exceeding exactly one threshold are Moderate, and those exceeding at least two thresholds are Challenging. Table 2 reports the proportion of each difficulty in the buffer at different PPO update steps.
@@ A.4 Ablation Study on the Number of Workers
We analyze the effect of the number of PPO workers, which determines how many agents collect trajectories in parallel. The only difference between our experimental setup and ACCEL (ACCEL) is the worker count: we use 16 workers for MiniGrid and 4 for BipedalWalker, whereas the original ACCEL paper uses 32 and 16, respectively. To ensure this configuration does not unfairly disadvantage TRACED, we additionally evaluate MiniGrid with 32 workers. Even under this setting, TRACED consistently outperforms ACCEL, indicating that our gains are robust to the PPO worker configuration (Table 7).
@@ Appendix G Overall Workflow
Algorithm 1 summarizes our overall UED procedure under the TRACED framework. We followed ACCEL (ACCEL) in all respects except the procedure used to compute task priority. At each iteration, with probability $1-p_{\mathrm{replay}}$ the teacher enters the exploration phase by sampling a fresh level $\theta\sim\mathcal{G}$, executing the current policy $\pi_{\phi}$ to collect a trajectory, and computing its approximated regret

### v5

@@ Abstract
Generalizing deep reinforcement learning agents to unseen environments remains a significant challenge. One promising solution is Unsupervised Environment Design (UED), a co‑evolutionary framework in which a teacher adaptively generates tasks with high learning potential, while a student learns a robust policy from this evolving curriculum. Existing UED methods typically measure learning potential via regret, the gap between optimal and current performance, approximated solely by value‑function loss. Building on these approaches, we introduce the transition-prediction error as an additional term in our regret approximation. To capture how training on one task affects performance on others, we further propose a lightweight metric called Co-Learnability. By combining these two measures, we present Transition‑aware Regret Approximation with Co‑learnability for Environment Design (TRACED). Empirical evaluations show that TRACED produces curricula that improve zero-shot generalization over strong baselines across multiple benchmarks. Ablation studies confirm that the transition-prediction error drives rapid complexity ramp‑up and that Co‑Learnability delivers additional gains when paired with the transition-prediction error. These results demonstrate how refined regret approximation and explicit modeling of task relationships can be leveraged for sample-efficient curriculum design in UED.
https://geonwoo.me/traced/
@@ 1 Introduction
Deep reinforcement learning (RL) has achieved remarkable success in games, continuous control, and robotics (Sutton et al., 1998). Ideally, we want agents that generalize robustly to a broad range of unseen environments. However, hand‑crafting a training distribution that captures all real‑world variability is intractable, and agents often overfit even large training sets, performing poorly out‑of‑distribution (Kirk et al., 2023; Korkmaz, 2024).
Unsupervised Environment Design (UED) tackles this by adapting the curriculum: a teacher module generates training tasks that challenge the student agent (Dennis et al., 2020). A popular class of UED methods measures task difficulty by regret, the difference between the optimal return and the agent’s achieved return, and uses this metric to guide curriculum design (Dennis et al., 2020; Jiang et al., 2021b; a; Parker-Holder et al., 2022; Azad et al., 2023a; Mediratta et al., 2023; Erlebach and Cook, 2024). Unfortunately, genuine regret requires knowing each environment’s optimal $Q^{*}$, which is infeasible in complex domains. Existing approaches, therefore, resort to coarse proxies such as Positive Value Loss (PVL) or maximum observed return (MaxMC) (Rutherford et al., 2024). In this paper, we refine regret estimation by augmenting PVL with a transition-prediction error term that captures how poorly a learned model predicts the environment’s dynamics. This combined signal provides a more faithful approximation of regret and a sharper basis for curriculum design.
We further introduce a metric called Co‑Learnability to quantify how training on one task benefits others. For instance, consider three 100‑word vocabulary tasks, Spanish, English, and Japanese, whose transfer patterns differ: because Spanish and English share many cognates, learning Spanish accelerates lexical access and boosts English accuracy (Ramírez et al., 2013; Costa et al., 2000; Lemhöfer and Dijkstra, 2004), reflecting high Co‑Learnability; in contrast, Japanese is typologically distant (Chiswick and Miller, 2005), so gains from Japanese may transfer unefficiently to English, reflecting low Co‑Learnability. We present a lightweight estimator of Co‑Learnability that leverages observed changes in approximated regret, avoiding any additional modeling overhead in the UED loop.
@@
We propose TRACED (Transition-aware Regret Approximation with Co-Learnability for Environment Design), which combines refined regret and Co-Learnability to yield a task-priority landscape (Figure 1). We evaluate TRACED on procedurally generated MiniGrid (MG) (Chevalier-Boisvert et al., 2023) and BipedalWalker (BW) (Romac et al., 2021). In MG, we compare against Domain Randomization (DR) (Jakobi, 1997), PLR (Jiang et al., 2021b), ADD (Chung et al., 2024), and ACCEL (Parker-Holder et al., 2022) (the strongest baseline). In BW, we additionally include the state-of-the-art (SOTA) method CENIE (Teoh et al., 2024).
@@ 2.1 Underspecified Partially Observable MDPs (UPOMDPs)
We model our environments as underspecified partially observable Markov decision processes (UPOMDPs) following Dennis et al. (2020). A UPOMDP is a tuple $\mathcal{M}=\langle A,O,\Theta,S,P_{0},P_{T},\mathcal{I},\mathcal{R},\gamma\rangle$, where $A$ is a finite set of actions, $S$ is a latent state space, and $O$ is an observation space. The observation function $\mathcal{I}:S\to\Delta(O)$ generates each observation $o_{t}$ given the true state $s_{t}$, the reward function $\mathcal{R}:S\times A\to\mathbb{R}$ and discount factor $\gamma\in[0,1)$ are shared across all levels. Crucially, $\Theta$ is a set of underspecified parameters that distinguish individual “levels”: for each $\theta\in\Theta$, the initial state is drawn from $P_{0}(\theta)\in\Delta(S)$ and transitions follow $P_{T}(s_{t+1}\mid s_{t},a_{t},\theta)$.
@@ 2.2 Unsupervised Environment Design (UED)
UED provides a series of levels with unknown parameters that are used to produce task environments automatically as a curriculum for the agent so as to efficiently train a single generalist policy $\pi_{\phi}$ across the entire parameter space $\Theta$ (Dennis et al., 2020). Recent UED methods maximize regret of the agent to generate a distribution of environments that guide effective learning (Jiang et al., 2021b; Parker-Holder et al., 2022; Teoh et al., 2024). The regret of policy $\pi$ is defined as the difference in expected reward between optimal policy $\pi^{*}$ and policy $\pi$, $Regret_{\theta}(\pi)=U_{\theta}(\pi^{*})-U_{\theta}(\pi)$.
Because the true optimal policy $\pi^{\star}$ is unknown in complex environments, existing methods approximate the instantaneous regret using proxy metrics. For example, PLR (Jiang et al., 2021a) evaluates two such proxies: Positive Value Loss (PVL) and Maximum Monte Carlo (MaxMC). PVL estimates the instantaneous regret as the average positive part of the Generalized Advantage Estimation (GAE)‐based Temporal Difference (TD) errors over an episode. PVL for an episode $\tau$ of length $T$ is defined as
@@
where $\gamma$ is the discount factor, $\lambda$ is the GAE coefficient and $\delta_{t}=r_{t}+\gamma\,V(s_{t+1})-V(s_{t})$ is the one‐step TD‐error at timestep $t$. MaxMC uses the highest undiscounted return observed on the task instead of a bootstrap target. Other criteria such as policy entropy, one‑step TD error, GAE, policy min‑margin, and policy least‑confidence have also been evaluated in curriculum learning contexts (Jiang et al., 2021b).
@@ 3.2 Task Priority Construction
A central challenge in UED is to decide which tasks to present to the agent at each step (Hughes et al., 2024). Intuitively, we want to (1) focus on tasks that remain challenging, while (2) exploiting tasks whose training yields transfer to others. To this end, we introduce two complementary quantities, Task Difficulty and Co‑Learnability, and combine them into a single Task Priority score. In the following sections, time $t$ indexes the teacher’s curriculum-selection cycles (not environment time).
@@
Definition 4 (Task Priority). We combine difficulty and Co-Learnability into a scalar score and then apply a rank transform, which replaces raw values with their relative order (e.g., the largest value receives rank 1, the next largest rank 2, etc.), as in ACCEL (Parker-Holder et al., 2022):
@@ 3.3 Overall UED Workflow
The overall UED algorithm follows the ACCEL loop (Parker-Holder et al., 2022), with the sole change that task scoring uses our Task Priority (Eq. 7) in place of Positive Value Loss (PVL) alone (Figure 3). At each curriculum update time $t$, with probability $d$ we sample a new level uniformly at random, and with probability 1d1-d we sample a level from the replay buffer according to its Task Priority score. Early in training, when the buffer is empty, we sample levels uniformly at random for a warm-up period. The warm-up length matches ACCEL’s setting.
@@ 4 Experiments
We evaluate TRACED on two procedurally generated domains: MiniGrid (MG) and BipedalWalker (BW). In both environments, we compare against DR, PLR, ACCEL (our primary baseline, since TRACED builds directly upon it), and ADD. In BW only, we additionally include the recent sota method CENIE (Teoh et al., 2024). While SFL (Rutherford et al., 2024) is not included in the main experimental suite, we provide a comprehensive comparison against SFL in the multi-agent JaxNav environment in Appendix Q, following its original evaluation protocol. Detailed descriptions of all baselines can be found in Appendix K.
For each domain, we track the emergent complexity of the sampled curricula during training and evaluate zero‐shot test performance on held‐out levels. To summarize test results, we report the median, interquartile mean (IQM), mean, and optimality gap using the rliable library (Agarwal et al., 2021). The optimality gap measures how far an algorithm falls short of a target performance level, beyond which further gains are deemed negligible. Accordingly, higher IQM and lower optimality gap values indicate better performance. Shaded regions in all plots denote 95% confidence intervals.
All methods use Proximal Policy Optimization (PPO) (Schulman et al., 2017) as the student agent. Following prior work (Parker-Holder et al., 2022; Teoh et al., 2024), we plot performance versus the number of PPO updates and report the corresponding environment interactions per update in Appendix F. Our only deviation from ACCEL is the number of PPO workers (i.e., parallel trajectory collectors), chosen for computational constraints: MiniGrid uses 16 workers instead of 32, and BipedalWalker uses 4 instead of 16. Consequently, some baseline scores may differ from those in the ACCEL paper. However, Appendix A.4 shows that TRACED has no worker-specific tuning and that reducing workers does not render the comparison unfair. We report TRACED at 10k updates (rather than 20k) because single-seed, long-horizon runs (45k PPO updates) on MiniGrid reveal post-convergence oscillations in both TRACED and ACCEL, and reporting earlier avoids this confound (Appendix B). Additional hyperparameters and architectures are provided in Appendix H.
@@ 4.1 Partially Observable Navigation
We evaluate our curriculum design on a partially observable maze navigation domain based on MiniGrid (Chevalier-Boisvert et al., 2023), as in prior UED work. Four representatives are shown in Figure 5. The agent observes a 147‑dimensional pixel observation and is trained for up to 20k PPO updates.
@@ 4.2 Walking in Challenging Terrain
We further validate our curriculum in the continuous‐control BipedalWalkerHardcore environment from OpenAI Gym (Brockman et al., 2016), as modified by Wang et al. (2019). This domain features a procedurally generated terrain controlled by eight parameters, terrain roughness, pit gap frequency, stump height, stair spacing, etc., that jointly determine locomotion difficulty. We consider the complete set of eight parameters in our design space. Figure 5 illustrates two representative terrains with varying stair heights and surface roughness. We evaluate zero‑shot transfer on six held‑out test terrains over 100 episodes each, averaged over five random seeds.
@@ 4.3 Ablation Study
To isolate component contributions, we compare TRACED at 10k updates with two ablations (Figure 11): ATPL only (TRACED - CL) and Co-Learnability only (TRACED - ATPL), each evaluated with five seeds. Results are reported following Agarwal et al. (2021). On MiniGrid, TRACED outperforms both variants across four metrics, indicating that each component contributes meaningfully to performance. The same trend holds on BipedalWalker (Appendix A). Ablations on the scaling factors $\alpha$ and $\beta$ further show that both elements play significant roles in TRACED and that carefully balancing ATPL and CL can yield additional gains (Appendix A.1). Detailed numerical results are provided in Appendix P.
@@ 4.4 Analysis on Curriculum Progression
To analyze how the curriculum evolves, we examine the BipedalWalker level buffer over time and categorize each generated level into three difficulty bands: Easy, Moderate, and Challenging, based on environment hyperparameters. Following Teoh et al. (2024), we set thresholds for Stump Height (2.4), Pit Gap (6), Ground Roughness (4.5), and Stairs Height (5). Levels exceeding no thresholds are labeled Easy, those exceeding exactly one threshold are Moderate, and those exceeding at least two thresholds are Challenging. Table 2 reports the proportion of each difficulty in the buffer at different PPO update steps.
@@ 5 Conclusion
Acknowledgments
This work was supported by IITP (RS-2024-00445087; 10%, RS-2025-25410841; 10%, No. 2019-0-01842; 10%), NRF (RS-2024-00451162; 20%, RS-2024-00454000; 20%, RS-2025-25419920; 20%), GIST (KH0870; 10%) funded by the Ministry of Science and ICT, Korea. Computing resource were supported by GIST SCENT-AI, AICA, KISTI, and NIPA.
@@ A.4 Ablation Study on the Number of Workers
We analyze the effect of the number of PPO workers, which determines how many agents collect trajectories in parallel. The only difference between our experimental setup and ACCEL (Parker-Holder et al., 2022) is the worker count: we use 16 workers for MiniGrid and 4 for BipedalWalker, whereas the original ACCEL paper uses 32 and 16, respectively. To ensure this configuration does not unfairly disadvantage TRACED, we additionally evaluate MiniGrid with 32 workers. Even under this setting, TRACED consistently outperforms ACCEL, indicating that our gains are robust to the PPO worker configuration (Table 7).
@@ Appendix G Overall Workflow
Algorithm 1 summarizes our overall UED procedure under the TRACED framework. We followed ACCEL (Parker-Holder et al., 2022) in all respects except the procedure used to compute task priority. At each iteration, with probability $1-p_{\mathrm{replay}}$ the teacher enters the exploration phase by sampling a fresh level $\theta\sim\mathcal{G}$, executing the current policy $\pi_{\phi}$ to collect a trajectory, and computing its approximated regret